OpenStudy (otonogold):

for the function f(x)= (3-4x)^2, find f^-1. determine whether f^-1 is a function. Can you show me how to do this ? Steps?

OpenStudy (dumbcow):

let f(x) be y then simply solve for "x" this will give you the inverse function

OpenStudy (otonogold):

f^-1(x) = [ (sq root x) -3 ] / (-4) << answer?

OpenStudy (dumbcow):

yes except the square root could be plus or minus which also means that f^-1 is NOT a function because for any positive x you have 2 "y" values

OpenStudy (otonogold):

So my answer would techinally be..? @dumbcow

OpenStudy (otonogold):

@jdoe0001 can you show me how to do this btw I cant see anything you draw because something wrong on my end of open study

OpenStudy (otonogold):

@agent0smith how is my answer incorrect or can you help me get a better understanding of what dumbcow said

OpenStudy (agent0smith):

y = (3-4x)^2 Solve that for x. Ugh... it's so hard to show w/o an equation editor. Remember when you take square root of both sides, you have to add a plus or minus sign on the left.

OpenStudy (anonymous):

it does not have an inverse

OpenStudy (agent0smith):

^Because of that plus or minus that you get when taking square roots.

OpenStudy (otonogold):

what do you mean by the plus or minus that I will get when taking square roots? especially in this f^-1(x) = [ (sq root x) -3 ] / (-4) as my answer..

OpenStudy (agent0smith):

When you take the square root of both sides (very first step), you must add a plus or minus sign.

OpenStudy (agent0smith):

I can't show it w/o the equation editor but if x^2 = 4 then x= 2 or x = -2 same thing here, except you just put a plus or minus sign in front of the sqrtx

OpenStudy (otonogold):

+/-?

OpenStudy (otonogold):

f ( f^( - 1 ) ( x ) ) = ( 3 - 4f^( - 1 ) (x ) )^2 x = ( 3 - 4f^( - 1 ) ( x ) )^2 +- sqrt ( x ) = 3 - 4f^( - 1 ) ( x ) - 3 + - sqrt ( x ) = - 4f^( - 1 ) ( x ) ( 3 / 4 ) + - ( 1 / 4 ) sqrt ( x ) = f^( - 1 ) ( x )

OpenStudy (otonogold):

right ?

OpenStudy (agent0smith):

I have no idea... this is far too hard to read like this but the last line looks right.

OpenStudy (otonogold):

basically x = (3 - 4y)^2 3-4y = +/- sqrtx -4y = - 3 +/- sqrtx y = 3/4 -/+ sqrtx

OpenStudy (agent0smith):

Yep.

OpenStudy (otonogold):

so this this is not a function unless the domain of x is restricted so that the sign on the square root is unambiguous?

OpenStudy (agent0smith):

^exactly :)